### v1

# Enrich the content of the image Using Context-Aware Copy Paste

Abstract
Data augmentation remains a widely utilized technique in deep learning, particularly in tasks such as image classification, semantic segmentation, and object detection. Among them, Copy-Paste is a simple yet effective method and gain great attention recently. However, existing Copy-Paste often overlook contextual relevance between source and target images, resulting in inconsistencies in generated outputs. To address this challenge, we propose a context-aware approach that integrates Bidirectional Latent Information Propagation (BLIP) for content extraction from source images. By matching extracted content information with category information, our method ensures cohesive integration of target objects using Segment Anything Model (SAM) and You Only Look Once (YOLO). This approach eliminates the need for manual annotation, offering an automated and user-friendly solution. Experimental evaluations across diverse datasets demonstrate the effectiveness of our method in enhancing data diversity and generating high-quality pseudo-images across various computer vision tasks.
Introduction
Deep Learning-based approach has become the major paradigm in many computer vision tasks, ranging from classification to segmentation. These approaches outperform traditional ones in terms of accuracy and generalization. However, the bottleneck of deep learning is the quality and quantity of the training dataset. To obtain a dataset, a large volume of images needs to be annotated, which is labour-intensive and time-consuming. In segmentation task, annotating a single image takes 1.5 hours(?). How to generate high-quality, highly realistic datasets has become a research hotspot in recent years.
Figure 1: Comparison between the Copy-Paste method (first row) and CACP (second row). The former overlooks the contextual relevance between the base and target images, leading to disharmony.
Previous data augmentation methods increase the diversity of images by applying operations such as flipping, rotating, and adding blur and noise. However, these techniques fail to enhance the content of images at the object level. To address this issue, the Copy-Paste method (?) was proposed. The idea is straightforward and intuitive: objects from target images are pasted onto source images at random positions, resulting in images with enriched content.
However, the Copy-Paste method has several drawbacks:
  • Context Neglect: The method overlooks the contextual relationship between the copied objects and the target images. For instance, a penguin is unlikely to appear in a desert, and a giraffe is unlikely to be found on a soccer field. These contextually inappropriate images contribute little to the practical significance of the augmented dataset.
  • Dependency on Masks: The original Copy-Paste approach relies on publicly available image-mask pairs to generate new images, requiring extra effort to extend its application boundaries. This process is not adaptable to scenarios where masks are unavailable.
To address the aforementioned issues, we propose a novel approach named Context-Aware Copy-Paste (CACP), leveraging large language models (LLMs). Our method integrates several NLP-based models to ensure contextual relevance between the source and target images. The main procedure is as follows: The BLIP model (?) is utilized to generate captions for the source images(the images to be augmented); Object365 (?), a dataset containing 365 distinct classes, serves as the target image set. For a given source image, we calculate a similarity score between its caption and the category names in Object365 using a BERT-based semantic similarity model. The category with the highest similarity score is selected, and an image from this category is randomly chosen as the target image; YOLO-obj365 is employed to detect objects in the target image. The bounding box of the detected object is then processed by SAM (?) to obtain a pixel-level mask;Finally, The object, guided by the mask, is pasted onto the source image.
Our approach can be applied to several computer vision tasks, including classification, object detection, and segmentation. Without requiring extra manual annotation, the target gallery can be easily extended to adapt to specific tasks. Moreover, the proposed module can be integrated with other techniques to enhance model performance. Our contributions can be summarized as follows:
  • We propose a data augmentation mechanism called Context-Aware Copy-Paste (CACP), which semantically bridges the source and target images. Additionally, this approach is easily extendable to custom tasks without requiring extra annotation.
  • We have demonstrated that robust segmentation results can be achieved by combining Grad-CAM and YOLO as prompt generators.
  • We evaluated our approach by conducting experiments across multiple tasks. The results indicate that our method outperforms previous data augmentation techniques.
Figure 2: Comparison between SAM segmentation results using different prompts. Single Bounding Box prompt(upper row)tend to get incomplete masks. Bounding Box with GradCAM points can generate more accurate and robust masks.
Related work
Copy-Paste
Crop-Paste has been widely used in semi-supervised scenarios for its efficiency.(??). Dovornik(?) first propose copy-paste approach for object detection task based on visual context, which boost the performance in VOC07(?) dataset.However, they only use voc2012 as their target gallery, which is hard to apply the proposed method to other specific scenarios.Besides, they use a CNN classifier to obtain context infomation, which is insufficient compared to our BLIP-based approach. Golnaz(?) first proposed Copy-Paste data augmentation method in Instance Segmentation. As described in their paper, They claim that simply pasting objects randomly is able to provide solid gain on top of baselines. Although their approach is easy to implement, the random-related approach generate images lose the concordance compared to real images. The distribution of co-occurrence of objects are ignored. Zhao(?) propose X-Paste, which leverage zero-shot recognition models CLIP to make the approach scalable. X-Paste provides impressive improvements over the CenterNet2. Viktor Olsson(?) propose ClassMix, which generate augmentations by mixing unlabeled samples by leveraging on the network’s predictions for respecting objects boundaries. We are inspired by their work to combine the vision language models with copy-paste to generate augmented images in a sufficient way.
vision language model and Zero-shot Segmentation
Vision-Language Pre-training (VLP) has recently made very encouraging breakthrough. The ability of zero-shot and image-text alignment makes it possible to assist the copy-paste pipeline.
CLIP (Contrastive Language-Image Pre-Training)(?) is a neural network trained on a variety of (image, text) pairs. It can be instructed in natural language to predict the most relevant text snippet, given an image, without directly optimizing for the task.
BLIP(?) is a new VLP framework which transfers flexibly to both vision-language understanding and generation tasks. It utilizes the noisy web data achieving SOTA results on several benchmarks, which also perform very well on zero-shot tasks.
SAM is a promptable segmentation model which is trained on the largest segmentation dataset so far(?). It can transfer zero-shot to new unseen distributions and tasks. The author claims that its zero-shot performance is competitive or even superior to prior fully supervised results. The only bottleneck of SAM is the prompt. The segmentation results may unstable or unintended given insufficient prompts. SAM is widely used to guide data augmentation. (?) introduce SAMAug, g, a novel visual point augmentation method for the SAM that enhances interactive image segmentation performance. However, the point they proposed lack semantic insight which is not efficient and meaningful.
Motivation
Although copy-paste data augmentation methods have significantly improved computer vision tasks, traditional crop-paste pipelines have two notable limitations. First, the crop-paste method is challenging to scale effectively. Second, the semantic gap between the source image and the target image remains unaddressed.
Scalability and Extendability
Previous copy-paste methods heavily rely on image-mask pairs to perform operations. However, preparing masks for images is costly and time-consuming, making it challenging to apply these methods in a generic manner. To address this issue, previous copy-paste approaches have resorted to using public datasets with pixel-level annotations such as VOC2017 and CamVid.
However, these datasets are limited in the number of categories they cover, thereby restricting the diversity of content in generated images. The table 1 provides descriptions of several public segmentation datasets (ADE20K (?), COCO (?), VOC2007 (?)). These datasets often cannot meet the specific requirements of scenarios. For instance, in a foreign object detection task, foreign objects may span hundreds of categories, making it impractical to rely on public datasets or manual annotation to prepare the dataset. An entirely automated crop-paste pipeline is needed to generate large quantities of high-quality images.
Dataset num Category resolution
VOC2007 9963 20 -
CamVid 701 32 480*360
CityScape 20000 30 2024*1048
coco 330k 80 640*480
ADE20K 25574 150 1650*2220
Table 1: Descriptions of public Segmentation datasets
Content dissonance
Previous work has often overlooked the issue of content relevance in the data augmentation process. Irrelevant objects are frequently pasted onto source images, providing minimal training benefit. As depicted in the figure below, images generated by traditional crop-paste methods contribute less to person-related computer vision tasks, as evidenced by corresponding Grad-CAM results.
We propose that incorporating highly relevant objects into source images at appropriate positions can enrich image content and expedite the training process. Experimental results validate our hypothesis: traditionally pasted parts fail to activate appropriately, while our method successfully triggers activations, enhancing the content relevant to desired classes.
Figure 3: GradCam comparison between the Copy-Paste(top row) and our context-aware copy paste(bottom row).
Method
In general, our CACP can be summarized in following parts: Gallery Preparation, Context-Aware, Crop Paste. In Gallery preparation part, target images are selected to provide object-level content enhancement; In context aware stage, source image and target image are bridged using BLIP and Bert-based similarity measurement tool. Once the preferred target image category is determined, object of interests in target image will be cropped and pasted onto the source image considering the size and position. The yolo and SAM will be leveraged to make the crop-paste process fully automatically. The details of each procedure will be described in follow paragraphs.
BLIP YOLO SAM total
Model Size(MB) 1824 14.2 2384 -
CPU(ms) 3687 97 3787 7571
GPU(ms) 1137 12 498 1647
Table 2: Model sizes and inference times in CPU and GPU
Figure 4: Our method’s pipeline involves leveraging BLIP and BERT to select the best-matched target image from a gallery. Subsequently, the corresponding mask is obtained using YOLO and SAM. A single base-target pair can generate multiple augmented images based on user preferences.
Reformulation
Given a source image sets $S_{S}$ and a target image $I_{T}$, our task is to find the most relevant image $I_{S}\in S_{S}$ and most relevant object $obj\in I_{S}$. Specifically, $I_{S}$ can be obtained as follow:
$I_{s}=\mathop{\arg\min\hskip 5.0pt}\limits_{I_{i}}\phi(I_{i},I_{T})$ (1)
where $\phi(\cdot)$ is the function to measure the semantic similarity between two images. Once $I_{S}$ is determined, $obj$ and corresponding mask $M$ can be inferred as follow:
$obj,M=\psi(I_{S})$ (2)
where $\psi(\cdot)$ are deep learning models, which take images as input and output coordinates(detection task) or labels of each pixel(segmentation task).
$I_{syn}=I_{S}\otimes M+I_{T}\otimes(1-M)$ (3)
where $I_{syn}$ is the generated image.
Data preparation
To enhance the diversity of our dataset, a substantial collection of images is essential for our gallery. In this study, we utilize Object365 (?) as our image gallery. Object365 encompasses 365 categories, featuring over 2 million images and 30 million bounding boxes. These images are characterized by high resolution and quality annotations. In contrast, COCO offers only 80 classes. Object365 significantly expands the range of target objects available for augmentation.
Additionally, we propose an alternative method to leverage images without bounding box annotations, thereby enhancing the applicability of our approach. Specifically, we assume that all images in the galleries are presented without bounding boxes or masks, and each image is labeled solely with its category name. This approach enables users to extend custom categories and adapt them to specific scenarios.
Context-Aware
Image Caption
To establish semantic coherence between the source and target images, it is crucial to recognize the contents of both images beforehand. Rather than solely detecting objects within the images, we employ a state-of-the-art Visual-Language pre-training (VLM) model as the content extractor. In this role, we utilize BLIP (Bootstrapped Language-Image Pre-training). Compared to object-detection methods, BLIP generates smooth and natural descriptions of input images, rather than isolated words. Furthermore, while object-level approaches may struggle to provide meaningful information when encountering unseen objects, BLIP consistently offers general information applicable to common scenarios.
Target object matching
In last step we obtain the caption of source image,namely $C(I_{src})$. Due to the large amount of the target image gallery, as a trade-off, we take the category name as the caption of the target image, annotated as $C(I_{T_{i}}(i=1,2,...,n))$,where n is the total number of categories. To determine the correlation between the $C(I_{src})$ and $C(I_{T_{i}})$, Bert-embedding is leveraged as our measurement tool. Below are samples using Bert-embedding to calculate the similarity in our work compared to traditional approaches.
Image caption category Bert
“Two teams are playing football games” soccer 0.94
pig 0.41
”A boy is dancing with a girl in the garden” person 0.89
goose 0.46
”A boy is standing near a red car” flower 0.51
truck 0.89
Table 3: Comparison between BERT-similarity and Cosine-similarity.
From the table we can find that Bert-based distance metric is preferred.
Crop-Paste
Once the category with the highest similarity score is determined, we randomly select an image from this category as the target image. The Segment Anything Model (SAM) is then employed as the pixel-level mask extractor. SAM is a single-shot segmentation model capable of segmenting any object based on prompts, such as bounding boxes or multiple points indicating the intended objects.
Prompts generation
To obtain these prompts, we utilize YOLO-365, a model trained on the Object365 dataset, to detect objects within the target images. For instance, if the YOLO-365 model detects a ’dog’ within a target image categorized under dogs, the bounding box of the dog is then forwarded to the SAM model along with the target image. SAM subsequently generates the corresponding segmentation mask for the detected object. Finally, guided by this generated mask, the pixels representing the dog are pasted onto the source images.
In our experimental trials, we observed that feeding pure bounding boxes into the Segment Anything Model (SAM) often results in unintended or incomplete masks for the corresponding target objects. To mitigate this issue, we propose an approach based on Grad-CAM (?) to achieve more accurate segmentation masks.
By inputting the target image into the Grad-CAM module, the resulting heatmap provides valuable positional information about the target. We then use this heatmap to sample points, which are combined with the bounding box as prompts for the SAM model. This hybrid approach, illustrated in Figure 2, improves the accuracy of segmentation results compared to using bounding boxes alone.
Scale and Position
To enhance the realism and harmony of the generated image, rescaling and rendering techniques are implemented. The cropped objects are rescaled according to a ratio interval based on our statistical analysis of the Object365 dataset. We traversal the images in Objects365, and record ratios of image pair, namely ¡obj1, obj2, ratio¿. Before pasting the object onto the source images, we extract object pair names and obtain the $ratio_{max}^{obj1,obj2}$ and $ratio_{min}^{obj1,obj2}$ from record.
Experiments
Methods Camvid
Car Pedestrian Building Road Sky Tree
U-Net(?) 0.776 0.447 0.764 0.872 0.872 0.834
U-Net+CACP 0.789 0.481 0.783 0.893 0.875 0.841
FPN(?) 0.792 0.432 0.783 0.897 0.884 0.867
FPN+CACP 0.813 0.479 0.797 0.903 0.885 0.881
PSPNet(?) 0.788 0.445 0.792 0.886 0.875 0.843
PSPNet+CACP 0.803 0.487 0.797 0.901 0.873 0.862
DeepLabV3(?) 0.792 0.461 0.803 0.908 0.891 0.875
DeepLabV3+CACP 0.811 0.493 0.812 0.922 0.887 0.893
DeepLabv3plus(?) 0.803 0.471 0.817 0.927 0.907 0.871
DeepLabv3plus+CACP 0.817 0.497 0.826 0.933 0.912 0.883
PAN(?) 0.794 0.501 0.806 0.931 0.883 0.868
PAN+CACP 0.812 0.513 0.821 0.937 0.891 0.891
Table 4: CACP works well across different segmentation architectures
Configuration
The experiment is conducted in pytorch platform. GPU is RTX 3090ti with 24GB memory. For classification task, Batch size is set to 16, loss function is cross entropy. Adam(?) is leveraged as optimizer, and learning rate is 0.001, epoch is set to 50; As for segmentation task, batch size is set to 8,loss function is dice loss,epoch is set to 20. As for detection task, we set batch size as 8, and epoch is set as 50. The models for segmentation task and object detection task we use the segmentation pytorch model and yolov5-s, respectively. Our segmentation models are implemented by segmentation models pytorch(?).
Metric
Precision is selected as the metric for classification task, which is calculated as below:
$Accuracy=\frac{Number\_correct\_predictions}{Totl\_number\_of\_Predictions}$ (4)
As for the segmentation task, mIoU is leveraged to evaluate the model’s performance.
$IoU=\frac{TP}{TP+FP+FN}$ (5)
$mIoU=\frac{1}{C}\sum_{c=1}^{C}IoU_{c}$ (6)
To evaluate the performance of the object detection model, we use mAP as out metric, which stands for Mean Average Precision. Here we set the threshold as 50%, which means IoU over 0.5 will be considered as
$mAP=\frac{\sum_{i=1}^{C}AP_{i}}{C}$ (7)
Evaluation Datasets
To comprehensively assess the viability of our proposed approach, we conducted experiments across three distinct datasets: Cat-Dog classification(?), and CityPersons(?). These datasets are representative of key tasks in the computer vision field: classification, segmentation, and object detection, respectively.
The Cat-Dog dataset contains 25,000 images, each labeled as either a cat or a dog. The CityPersons(?) dataset is a subset of Cityscapes, focusing solely on person annotations. It includes 2,975 images for training and 500 images for validation. The Cambridge driving Labeled Video Database(CamVid)(?) is the first collection of videos with object class semantic labels, complete with metadata. The dataset contain over 700 images with pixel-level annotation. The annotation of images cover 32 class labels, include building, car, tree, sky, etc.
Methods Cat-Dog CityPersons
(Acc) (mIoU) (mAP)
B 0.927 0.914 0.557
B+CP 0.941 0.897 0.561
B+aug 0.957 0.903 0.567
B+aug+cp 0.962 0.911 0.571
B+CACP 0.969 0.929 0.577
B+CACP+aug 0.974 0.938 0.591
Table 5: Results between copy-paste(CP) and context-aware Copy- Paste(CACP) in classification, segmentation and detection tasks.
Results
Across initialization
To valid the robustness of CACP across different initialization, we conduct experiments on CamVid based on two different initialization configurations, namely imagenet(?) pretrained and normal initialization. As illustrated in Fig .5, the results with CACP outperform the one without CACP in both configurations: 0.803 to 0.817 in random parameter initialization; 0.842 to 0.864 in imagenet pre-trained config.
Figure 5: Comparison between different initialization configuration W/Wo CACP.
Across Tasks
We conduct experiments on different computer vision tasks to valid the usage scenarios of CACP, namely image classification tasks on Cat-Dog, image segmentation and object detection on CityPersons. The item in column Methods indicate different combinations. B indicates base model. In classification task, which is Resnet-50, in segmentation is DeepLabv3, and in object detection is YOLOv5-s. CP and CACP indicate random crop-paste and context-aware copy-paste, respectively. aug is a combination of traditional data augmentation techniques, including flip, colorjitting, random noise, which is provided by lib Albumentation. As illustrated in Table 5, all augmentation techniques have the ability to boost the model. CACP outperforms CP and aug across all three tasks.
Across Architectures
To valid the effectiveness of our method across different architectures, we conduct experiments on CamVid(?) dataset using popular encoder-decoder architectures, namely U-net(?), FPN(?), PSPNet(?), DeepLabv3(?), DeepLabV3+(?), and PAN(?). The experiment is conducted with or without CACP operation. To observe the effect of different categories, we select six classes: [car,pedestrian,building,tree,sky,road].
Across Partition
To valid the effect of the number of augmented images, we conduct the experiments on CamVid. 1/n indicates 1/n of training images have been augmented using CACP. It is important to note that the total number of training image is fixed, only the ration of augmented and non-augmented varies. The result illustrates that the performance increases with the rise of partition from 1/8 to 1/2, the trend is stable in all 4 experiments. The increase is saturated when partition is over 1/2.
Speed up convergence
We have noticed that CACP contributes to speed up the training process. As illustrated in Fig .6, the CACP augmented one converges rapidly compared to the wo-CACP one. The loss is stable around epoch 15, while the wo-CACP is still not fully converged after epoch 19.
Figure 6: dice loss during training with or without CACP.
Effect of GradCAM Prompt
To improve the robustness and stability of the output masks of SAM, we propose the GradCAM-guided approach to generate points+bounding box prompts. During our trials, we notice that how many points in GradCAM to select affect the final segmentation mask. To determine the best point numbers, we conduct the following experiment on CamVid. random indicates the point is sampled randomly inside bounding box. CAM(n) means extract n points in GradCAM map with high value. The Table7 indicates that extra prompts can improve the accuracy of the SAM; CAM-based point prompts are better than random points. The prefered number of points is around 3 to 5.
Methods DeepLabv3
1/8 1/4 1/2 Full
PSPNet 0.872 0.891 0.887 0.893
U-NET 0.851 0.873 0.879 0.877
PAN 0.862 0.871 0.877 0.874
DeepLabv3 plus 0.883 0.903 0.901 0.907
Table 6: Results between copy-paste(CP) and context-aware Copy- Paste(CACP) in classification, segmentation and detection tasks.
bbox +rand(1) +CAM(1)+ CAM(3) +CAM(5)
0.734 0.841 0.927 0.934 0.933
Table 7: mIoU across different prompts.
Methods Camvid
car pedestrian building tree traffic lane traffic cone
SAM+B
SAM+B
SAM+B+random(1)
SAM+b+CAM(1)
SAM+b+CAM(3)
SAM+b+CAM(5)
Table 8: Caption
Discussion
n this paper, we propose a context-aware copy-paste (CACP) data augmentation approach that serves as a versatile plug-in module for various computer vision tasks, eliminating the need for additional manual annotation. CACP proves highly efficient for custom segmentation tasks in specific scenarios. Instead of annotating thousands of pixel-level masks, fewer than 100 annotated base images are sufficient to generate a multitude of highly realistic pseudo pixel-level images for training models. This process is both time-saving and scalable, allowing users to customize their target gallery according to specific task requirements.
As a case study, we applied CACP to our Railway obstacle detection project (?). With just a single annotation mask (due to fixed camera placement), we trained a universal obstacle detection model that outperformed previous object detection-based approaches. CACP proves particularly effective in semi-supervised segmentation by generating high-quality pseudo images. Moving forward, we aim to develop more user-friendly tools to enhance accessibility and extendability.